a fishpond has a radius of 60cms and is 40cm deep. The ponds needs refilling. a garden hose can fill it at the rate of 2 litres per minute, how long will it take the pond to fill.
thanks

Answer by Theo(2967)

1 liter = 1000 cubic centimeters.

The radius of the pond surface is 60 centimeters.

The pond is 40 centimeters deep.

The volume of the pond would be the area of the pond surface times the depth.

The area of the pond surface is equal to pi*r^2 = pi*60^2 = pi*3600 = 11309.73355 square centimeters.

The volume of the pond is equal to the area of the pond surface * 40 = 452389.3421 cubic centimeters.

The number of cubic centimeters per liter is 1000.

The number of liters in the pond is equal to 452389.3421 cubic centimeters / 1000 cubic centimeters per liter = 452.3893421 liters.

At the rate of 2 liters per minute, it will take 452.3893421 / 2 = 226.1946711 minutes to refill the pond.

That is equivalent to 3 hours and 46.19467106 minutes.